OpenStudy (anonymous):

2n^4 * 6n^4 Do i add the bases or multiply them?

OpenStudy (rulnick):

12 n^8

OpenStudy (rulnick):

neither. 2 n^4 6 n^4 = 2 (6) n^4 n^4 = 12 n^8

OpenStudy (phi):

if you forget how to multiply n^4 * n^4 remember what n^4 means: n*n*n*n so n^4 * n^4 = n*n*n*n * n*n*n*n = n^8 it's definitely easier to add the exponents 4+4=8

OpenStudy (anonymous):

but what do i do about the 2n and 6n?

OpenStudy (phi):

2*6 =12 n*n= n^2 12n^2

OpenStudy (rulnick):

Is it (2n)^4 or 2n^4? Is the 2 being raised to the 4th power? If not, there's nothing else to do. The 2 and 6 can be multiplied to get 12.

OpenStudy (phi):

when you multiply, you can re-arrange things: 2n*6n= 2*6*n*n= 12n^2

OpenStudy (rulnick):

If it's (2n)^4 (6n)^4 then the answer is instead 20736 n^8.

OpenStudy (anonymous):

ok thanks and can you help me do 7v^3*10u^3v^5*8uv^3

OpenStudy (phi):

just to be clear, when people write 7v^3 (for example) that means 7*v*v*v or 7*v^3 so, first sort things: all the numbers times each other, all the bases times each other (do it alphabetical)

OpenStudy (phi):

7v^3*10u^3v^5*8uv^3 = 7*10*8 * (u^3 *u)* (v^3 * v^5 * v^3)

OpenStudy (phi):

Can you finish?

OpenStudy (anonymous):

i dont kow man

OpenStudy (phi):

7v^3*10u^3v^5*8uv^3 = 7*10*8 * (u^3 *u)* (v^3 * v^5 * v^3)

OpenStudy (phi):

you can do it step by step 7*10*8= u^4*u = v^3 * v^5 *v^3 = then put it together.